5. Calculate the z-transform of x[n] = 0.2u[n – 3], and determine its region of convergence.
Answer #1

Solution : siven signal is- 20 [m] = (0-2)” [-30 a [m 3] => ر n<3 n 23 x En] = (0.2). n23 { n23 = X (2) Z[3] = we know that- X(2) = (0.008 23) I-O o.gz (0) 0.008 X (2) = 23 (1-0-22) → for Region of convergence. → (2+0) o and | 0.22) <1 (or) eauiv